OpenStudy (anonymous):

Need help quickly please! The position of an object at time t is given by s(t) = -8 - 9t. Find the instantaneous velocity at t = 1 by finding the derivative.

OpenStudy (ipwnbunnies):

The derivative of a position function is the velocity function. Can you find the velocity function?

OpenStudy (anonymous):

I have no idea how. Could you take me through it?

OpenStudy (ipwnbunnies):

Oh man, I'm not good at teaching derivatives.

OpenStudy (ipwnbunnies):

We can take it one term at a time. So we need to find the derivative of f(x). We can find the derivative of each term.

OpenStudy (anonymous):

Alright so what do i do?

OpenStudy (ipwnbunnies):

Let's do -8 first. The derivative of a constant is ALWAYS ZERO. Aight? We have this so far: f(x) = -8 - 9t f'(x) = v(t) = 0 + ____

OpenStudy (anonymous):

ok ya that makes sence

OpenStudy (ipwnbunnies):

I shouldn't have put that plus sign, but it's ok. We can look at -9t as one term now.

OpenStudy (anonymous):

sense*

OpenStudy (ipwnbunnies):

With -9t, we use the Power Rule for derivatives: \[\frac{d}{dx}x^n = n*x^{n-1}\]

OpenStudy (anonymous):

hmm ok how do i use it

OpenStudy (ipwnbunnies):

Plus, we can isolate the -9 as a multiple of t, which is another property of derivatives. It can be written as this: -9 * t^1, right?

OpenStudy (ipwnbunnies):

Can you apply the power rule to this term to find its derivative?

OpenStudy (anonymous):

d/d(9) 9^n=n*9^n-1?

OpenStudy (ipwnbunnies):

Uh no, you're apply the rule to t. I'll show you now. We also keep -9 as a multiple: \[\frac{d}{dt} -9t = -9*\frac{d}{dt}t^1 = -9*1*t^{0} = -9 \]

OpenStudy (ipwnbunnies):

You see how I applied the rule? The third step?

OpenStudy (anonymous):

ya i think so... what's next

OpenStudy (ipwnbunnies):

Well, that's the last term of the function. We write it nicely now: f(x) = -8 - 9t v(t) = 0 - 9 = -9

OpenStudy (anonymous):

so the answer is -9? wow thanks so much!

OpenStudy (ipwnbunnies):

v(t) = -9; When you plug in anything for t, what value will you get?

OpenStudy (anonymous):

depends on the value of t?

OpenStudy (ipwnbunnies):

The velocity function looks like a horizontal line through y-value -9. So, at any time 't', the velocity will be -9 units

OpenStudy (ipwnbunnies):

Yeah, it depends on the value of t. Not always will we get a constant for the velocity function.

OpenStudy (anonymous):

oh ok, thanks so much :)